# Dream-IF: Dynamic Relative EnhAnceMent for Image Fusion

### v1

@@ Abstract
Image fusion aims to integrate comprehensive information from images acquired through multiple sources. However, images captured by diverse sensors often encounter various degradations that can negatively affect fusion quality. Traditional fusion methods generally treat image enhancement and fusion as separate processes, overlooking the inherent correlation between them; notably, the dominant regions in one modality of a fused image often indicate areas where the other modality might benefit from enhancement. Inspired by this observation, we introduce the concept of dominant regions for image enhancement and present a Dynamic Relative EnhAnceMent framework for Image Fusion (Dream-IF). This framework quantifies the relative dominance of each modality across different layers and leverages this information to facilitate reciprocal cross-modal enhancement. By integrating the relative dominance derived from image fusion, our approach supports not only image restoration but also a broader range of image enhancement applications. Furthermore, we employ prompt-based encoding to capture degradation-specific details, which dynamically steer the restoration process and promote coordinated enhancement in both multi-modal image fusion and image enhancement scenarios. Extensive experimental results demonstrate that Dream-IF consistently outperforms its counterparts.
1 Introduction
Image fusion aims to integrate essential information from multi-source images captured by various sensors, producing a single comprehensive image. Multi-modal (visible v.s. infrared) image fusion [43, 20, 15] has been used in a wide range of applications, such as auto driving [10], unmanned aerial vehicles [12], forest fire monitoring [18], etc. Images captured by different imaging sensors often have different characteristics. The infrared images intrinsically avoid visible-light interference but often lack fine texture details. In contrast, the visible images capture abundant color and details in well-lit areas, while may suffer significant quality degradation in complex environments. The fused image preserves the benefits of visible and infrared images while minimizing their limitations by utilizing the complementary information from the respective modalities.
However, due to sensor malfunctions or environmental disturbances, noise is often present in real-world visible and infrared images [23, 26] resulting in degradation and low quality. Some existing methods [35, 2] enhance the low-quality images by learning the distribution of datasets and have achieved impressive results. More recently, some approaches [3, 28, 33] effectively handle degradations by capturing long-range pixel interactions through multi-head attention and feed-forward networks, leveraging the relationship to model the restoration process.
@@ 1 Introduction
Intuitively, from the perspective of multi-modal complementarity, image fusion acquires the dominant information of different modalities, which reversely indicates the relative non-dominant in the other modality, specifically tailoring regions that should be enhanced regardless of degradation issues or not.By leveraging the relative dominance in image fusion, the difficulty of image degradation recovery can be mitigated. Simultaneously, the use of image degradation tasks can further enhance the adaptability of image fusion to various degradation scenarios.
Based on this motivation, we propose a Dynamic Relative EnhAnceMent framework for Image Fusion (Dream-IF), jointly boosting the performance of image fusion and enhancement. Unlike most existing methods, Dream-IF first introduces the natural multi-modal complementarity from image fusion to enhance cross-modal image quality, especially for the corresponding non-dominant regions of the respective modality.
To perform relative dominance-aware enhancement, we designed the relative enhancement (RE) module to dynamically capture the relative dominance of each modality, which is subsequently employed to facilitate cross-modal enhancement of the relatively weak-quality regions. Consequently, we can obtain the enhanced multi-modal features. By revealing the comprehensive nature of image fusion and converting the dominant regions of one modality to guide the enhancement of the other modality, we enhance the model’s capacity for image fusion in complex conditions, making the fusion model more robust to degradations. To the best of our knowledge, this work is the first to explore the intrinsic correlation of image enhancement and fusion using comprehensive relative dominance, rather than treating them as separate data-driven tasks or merely integrating them into a unified framework. Our model copes with robust image fusion for image degradation and even broader image enhancement. The main contributions can be summarized as follows:
  • We for the first time investigate the complementarity of image fusion from the perspective of image enhancement. Based on this, we propose a Dynamic Relative EnhAnceMent framework for Image Fusion, termed Dream-IF, which is qualified to perform robust image fusion for image degradation and even broader enhancement.
@@
2 Related work
Image Fusion. Image fusion focuses on extracting effective information from multi-source images and producing an image containing complementary information. Traditional methods often perform image fusion by mathematical transformations, with multiscale analysis [41], sparse representation [39], subspace learning [8], etc. With the development of deep learning, CNN-based methods have made significant advancements. DenseFuse [15] pioneered the use of deep learning models for fusing infrared and visible images. Inspired by transform-domain image fusion, IFCNN [40] introduced a general image fusion framework for various fusion tasks. Furthermore, U2Fusion [31] addresses various fusion tasks through a unified, unsupervised image fusion network. Similar fusion tasks with aligned objectives can facilitate the integration of complementary information through cross-task commonality. More recently, TC-MoA [45] dynamically selects task-customized mixture of adapters to capture task generality while preserving unique task characteristics. TTD [1] proposed an effective test-time fusion paradigm based on generalization theory proving dynamic fusion is superior to static fusion. Considering the potential low-quality of different sensors, some researchers [6, 7] have begun to recognize the correlation between image restoration and fusion. DIVFusion [27] enhances low-light images to produce a daytime-like fused image with improved visual perception. Different from these techniques, we first attempt to introduce the naturally complementary dominance of image fusion to dynamically enhance the image quality of non-dominant modality, improving the overall fusion results.
Image Restoration. In recent years, image restoration techniques [16, 35, 3, 28] have achieved significant progress. Particularly, blind restoration [29, 11, 25] has gained considerable attention. Blind degradation problem is inverse where the degradation is not explicitly known, i.e., non-blind methods can model degradation using prior knowledge (e.g., a predefined blur convolution kernel) while blind methods only assume knowledge of the type of degradation (e.g., blurring or JPEG compression)[4]. ECycleGAN  [29] tackled the blind reserve problem as an image-to-image task while preserving the fidelity of the reconstructed image. BlindDPS [5] solved blind inverse problems by constructing a separate diffusion prior for the forward operator. AirNet [14] leveraged consistency to learn the degradation representation, recovering various degraded conditions. PromptIR [24] introduces a prompt-based learning approach to encode degradation-specific features with prompts and dynamically guide all-in-one image restoration. Despite the extensive exploration of image restoration, these approaches fail to consider the inherent correlation between image restoration and fusion, specifically using the complementarity in image fusion to assist restoration.
@@ 2 Related work
Figure 2: An overview of the proposed Dream-IF. We introduce the Relative Enhancement block, which implicitly enhances non-dominance representations by leveraging relative information during the fusion and restoration process. This module captures the relative dominance inherent in the complementary nature of the image fusion model and uses it to facilitate both cross and self enhancement, ultimately producing the enhanced feature.
3 Method
@@
We present the pipeline of our Dream-IF in Fig. 2, a dynamic relative enhancement framework acting on the fusion model, which leverages relative dominant prompts to perform cross-modal enhancement for non-dominant regions, thereby facilitating a more robust fusion. First, we describe the Dream-IF fusion pipeline in Sec. 3.1. Next we illustrate the primary component of our work, Relative Enhancement in Sec. 3.2. Finally, we clarify the loss function of Dream-IF in Sec. 3.3
3.1 The Overview of Dream-IF
@@
Dream-IF consists of two sets of encoders and decoders containing Restormer blocks [35] corresponding to different modalities, which can be defined as ${\{\mathcal{E}^{m},\mathcal{D}^{m}|m\in\{i,v\}\}}$, where $\mathcal{E}$ and $\mathcal{D}$ represent the encoder and decoder, respectively. Given the visible image $I^{v}\in\mathbb{R}^{H\times\ W\times 3}$ and the infrared image $I^{i}\in\mathbb{R}^{H\times\ W\times 3}$, the feature of each modality can be obtained as follows:
$F^{m}=\mathcal{D}^{m}(\mathcal{E}^{m}(I^{m})),$ (1)
@@
where $F^{m}(m\in\{i,v\})$ represents the latent feature of the given modality and the $F^{m}_{l}$th denote the feature at the $l$th. Our goal is to fuse the complementary information from the two source images and obtain a fused image $I^{f}$. Thus, the fusion module can be formulated as $I^{f}=\mathcal{F}(\sum_{k\in m}{(F^{k})})$, where consists of a Restormer block and a convolution layer to reconstruct the fused image from features.
3.2 Relative Enhancement
@@
The encoder and decoder tend to extract contemporary information from source images, which means that when a region of one modality is advantageous, another modality is probably disadvantageous. Inspired by the comprehensive nature of image fusion, we proposed a relative enhancement module, to enhance the deficient regions of one modality by capturing the relative dominance of another modality, thus generating a more robust image enhancement and fusion.
Relative Dominance. As discussed above, the image fusion process inherently aims to exploit complementary information. We define the fusion weight for each source feature, obtained by the Refine block, which consists of a Restormer block and a convolution, as the relative dominance (RD), where $RD^{m}=\sigma Conv(F^{m})$, and $\sigma$ is an activation function. It captures the dominance of each source in the fusion process and leverages the dominant modality to restore non-dominant areas.
@@
Relative Enhancement. In image fusion, the dominance of one modality can be approximated by the non-dominance of another modality, which should be appropriately enhanced. Thus, the primary goal of the relative enhancement (CE) module is to boost the representation of non-dominant feature by facilitating the relative dominance of another modality.
@@
To naturally integrate the $RD$ and the feature to be strengthened, we present the cross Enhancement and self Enhancement module to enhance. The $RD$ denotes the relative dominance from the other modality and indicates its deficient parts to be enhanced, thereby efficiently guiding the reconstruction process. Specifically, the cross Enhancement (CE) module addresses deficiencies in one modality by complementing it with information from the other modality. It leverages the complementary nature of the modalities, filling in the gaps of the defective feature with the more dominant or reliable feature from the other modality. This ensures a more complete and accurate representation of the scene. The restoration process is guided by the dominant region in the other modality, where the strength of one modality helps identify areas in the other that require enhancement. By focusing on these dominant regions, the module effectively restores the missing or degraded information in the subordinate modality, resulting in a more coherent and refined output. It can be formulated as
@@
Further to enhance itself and restore degeneration, to make our model fundamentally resistant to degradation [24], we generate a Self Enhancement (SE). This is calculated as:
$\hat{F}^{m}=F^{m}+\textbf{SE}(\textbf{RP}(F^{m},RD,z),F^{m})$ (3)
where $z$ is a set of learnable parameters and RP is the Relative Prompt for restoration, which dynamically generates enhancement prompts based on relative dominance and the input feature. The Restormer block in RP operates along the channel dimension and has linear complexity, suppressing less informative features while allowing only the useful ones to propagate through the network.
Afterward, the refined feature is sent to the next block in the decoder, where it is transformed into the denoised and restored feature:
@@
where $\widetilde{F}=Conv(\textbf{Cat}[\dot{F};\hat{F}])$. This enhanced feature is then passed to the fusion block, which combines the features from the two modalities and generates the final fused image.
3.3 Loss Function
@@
Specifically, the pixel loss is denoted as:
$\mathcal{L}_{pixel}=\frac{1}{HW}||I^{f}-max(I^{v},I^{i})||_{1}.$ (6)
The gradient loss aims to preserve texture details in the fused image by leveraging information from both modalities. It can be formulated as:
$\mathcal{L}_{grad}=\frac{1}{HW}||\nabla I^{f}-max(\nabla I^{v},\nabla I^{i})||_{1},$ (7)
where $\nabla$ represents the Sobel operator, which extracts texture information from the image. The SSIM loss maintains the structural similarity between the reconstructed image and the original image. It can be expressed as:
$\mathcal{L}_{SSIM}=(1-SSIM(I^{f},I^{v}))+(1-SSIM(I^{f},I^{v}).$ (8)
The color loss preserves color by calculating the similarity of the CbCr channels in the YCbCr color space. It can be expressed as:
$\mathcal{L}_{color}=\frac{1}{HW}||CbCr(I^{f})-CbCr(I^{v}))||_{1}.$ (9)
4 Experiments
@@
4.1 Experimental Setting
Implementation Details. Our experimental framework is implemented using PyTorch [21] and executed on an NVIDIA 3090 GPU. The proposed architecture adopts a U-Net-based design with a symmetric 4-level encoder-decoder structure, where the encoder extracts hierarchical features and the decoder reconstructs the output image. Each level of the encoder and decoder incorporates multiple Restormer [35] blocks, with the number of blocks progressively increasing from higher to lower levels. The complete architectural details, including layer configurations and hyperparameters, are provided in the supplementary material for reproducibility.
Table 1: Quantitative comparison with SOTAs without degradation. The red/blue indicates the best and runner-up.
@@ 4.1 Experimental Setting
EI $\uparrow$ AG $\uparrow$ PSNR $\uparrow$ $Q^{AB/F}$uparrow VIFF $\uparrow$ EI $\uparrow$ AG $\uparrow$ PSNR $\uparrow$ $Q^{AB/F}$uparrow VIFF $\uparrow$ EI $\uparrow$ AG $\uparrow$ PSNR $\uparrow$ $Q^{AB/F}$uparrow VIFF $\uparrow$
DIDFuse (IJCAI’20) $9.06$ $3.77$ $30.66$ $0.27$ $0.35$ $5.95$ $2.24$ $37.42$ $0.27$ $0.49$ $8.15$ $3.13$ 31.591.591.591.59 $0.43$ $0.62$
IFCNN (InF’20) $12.52$ $5.09$ $31.52$ $0.48$ $0.63$ $9.32$ $3.56$ $37.72$ $0.55$ $0.66$ $12.24$ $4.74$ $31.49$ $0.62$ $0.62$
STDFusion (TIM’21) $9.83$ $4.09$ $32.22$ $0.35$ $0.38$ $10.07$ $3.78$ $37.76$ $0.585$ $0.72$ $11.87$ $4.61$ $32.58$ $0.59$ $0.75$
TarDAL (CVPR’22) $11.90$ $4.92$ $28.68$ $0.41$ $0.54$ $6.32$ $2.41$ $35.16$ $0.26$ $0.41$ $10.44$ $4.33$ $23.79$ $0.29$ $0.49$
CDDFuse (CVPR’23) $13.31$ $5.40$ $32.13$ $0.58$ $0.66$ $10.07$ $3.78$ $37.76$ $0.59$ $0.72$ $12.45$ $4.80$ $30.75$ $0.52$ $0.58$
EMMA (CVPR’24) $14.08$ $5.56$ 32.96 $0.55$ $0.65$ $10.13$ $3.77$ $38.09$ $0.62$ $0.74$ 14.00 5.08 33.48 $0.59$ $0.76$
Text-IF(CVPR’24) 15.01 6.02 $31.97$ 0.70 0.85 $10.16$ $3.80$ 42.59 0.62 0.87 $13.35$ $5.03$ $29.26$ 0.64 0.79
TTD (NeuIPS’24) $14.01$ $5.51$ $31.29$ $0.63$ $0.73$ $9.71$ $3.71$ $37.60$ $0.55$ $0.66$ $12.98$ $5.02$ $31.47$ $0.60$ $0.64$
Text-Dif(NeurIPS’24) $12.56$ $4.85$ $26.07$ $0.40$ $0.52$ 10.36 3.84 $25.05$ $0.44$ $0.70$ $7.51$ $2.81$ $28.72$ $0.15$ $0.16$
Dream-IF 15.17 6.07 32.33 0.73 0.86 10.51 3.94 41.17 0.69 0.91 13.37 5.13 33.21 0.69 0.82
Datasets. Our experiments are conducted on three widely recognized publicly available datasets: LLVIP, MSRS, and M3FD. Further details about the datasets are provided in the supplement. Following [17], utilizing $12,025$ image pairs from LLVIP dataset for training and $70$ image pairs for testing. To rigorously assess the generalization capability of the proposed method, we tested the model on both the MSRS and M3FD datasets without any fine-tuning.
Competing Methods. We compared our approach with seven recent competing methods, including DIDFuse [42], IFCNN [40], STDFusion [19], TarDAL [17], CDDFuse [43], EMMA [44], Text-IF [34], TTD [1] and Text-DiFuse [36].
Evaluation Metrics. We evaluated the fusion results quantitatively on five metrics, including the edge intensity (EI), average gradient (AG), Peak signal-to-nosie ration (PSNR), gradient-based similarity measurement $Q^{AB/F}$ [22], and visual information fidelity for fusion (VIFF) [9].
4.2 Comparison without Degradation
@@
Qualitative Comparison. Our comprehensive evaluation reveals that while all competing fusion methods are capable of combining primary features from infrared and visible images to some extent, our proposed method exhibits two distinct advantages that significantly enhance fusion quality:
@@
First, our method excels in emphasizing the critical characteristics of infrared images while preserving complementary visible information. As illustrated in Fig. 3 (LLVIP dataset, red boxes), the visible images in low-light conditions are often blurred, whereas the infrared images provide essential complementary details. Competing methods, however, suffer from significant infrared content loss in the fusion results, as evidenced by the dark and indistinct details of the cars in the green boxes. Although CDDFuse shows competitive visual performance in the red boxes, it fails to retain sufficient infrared information, leading to suboptimal fusion in the green boxes. In contrast, our method enhances vulnerable modalities through a robust feature integration mechanism, effectively preserving both infrared and visible information. This results in superior perceptual quality and improved information retention, as demonstrated by the clear and detailed fusion results in both highlighted regions.
Second, our method achieves superior preservation of visible contrast and information, particularly in challenging scenarios involving extreme lighting conditions (e.g., overly dark or overexposed images). This is prominently illustrated in the MSRS dataset, where the green boxes highlight dark regions and the red boxes indicate overexposed areas. While other methods struggle to balance the fusion of both modalities under such conditions, our method demonstrates remarkable robustness. It effectively integrates visible and infrared information, leveraging their inherent complementarity to compensate for weaknesses in individual modalities. For instance, in the dark regions (green boxes), our method retains rich visible details by incorporating infrared thermal information, while in overexposed regions (red boxes), it suppresses over-saturation and preserves structural integrity.
Figure 3: Qualitative comparisons of various methods on representative images selected from the LLVIP, MSRS and M3FD dataset.

### v2

@@ Abstract
Image fusion aims to integrate comprehensive information from images acquired through multiple sources. However, images captured by diverse sensors often encounter various degradations that can negatively affect fusion quality. Traditional fusion methods generally treat image enhancement and fusion as separate processes, overlooking the inherent correlation between them; notably, the dominant regions in one modality of a fused image often indicate areas where the other modality might benefit from enhancement. Inspired by this observation, we introduce the concept of dominant regions for image enhancement and present a Dynamic Relative EnhAnceMent framework for Image Fusion (Dream-IF). This framework quantifies the relative dominance of each modality across different layers and leverages this information to facilitate reciprocal cross-modal enhancement. By integrating the relative dominance derived from image fusion, our approach supports not only image restoration but also a broader range of image enhancement applications. Furthermore, we employ prompt-based encoding to capture degradation-specific details, which dynamically steer the restoration process and promote coordinated enhancement in both multi-modal image fusion and image enhancement scenarios. Extensive experimental results demonstrate that Dream-IF consistently outperforms its counterparts. The code is publicly available. [^1] https://github.com/jehovahxu/Dream-IF
Introduction
Image fusion aims to integrate essential information from multi-source images captured by various sensors, producing a single comprehensive image. Multi-modal (visible v.s. infrared) image fusion (Zhao et al. 2023; Ma et al. 2022; Li and Wu 2018) has been used in a wide range of applications, such as auto driving (Huang et al. 2022), unmanned aerial vehicles (Jasiunas et al. 2002), forest fire monitoring (Liu et al. 2023), etc. Images captured by different imaging sensors often have different characteristics. The infrared images intrinsically avoid visible-light interference but often lack fine texture details. In contrast, the visible images capture abundant color and details in well-lit areas, while may suffer significant quality degradation in complex environments. The fused image preserves the benefits of visible and infrared images while minimizing their limitations by utilizing the complementary information from the respective modalities (Ma, Ma, and Li 2019; Liu et al. 2024).
@@ Introduction
However, due to sensor malfunctions or environmental disturbances, noise is often present in real-world visible and infrared images (Plotz and Roth 2017; Tang, Liu, and Su 2014) resulting in degradation and low quality. Some existing methods (Zamir et al. 2022; Chen et al. 2021) enhance the low-quality images by learning the distribution of datasets and have achieved impressive results. More recently, some approaches (Chen et al. 2022; Wang et al. 2022; Yang et al. 2023) effectively handle degradations by capturing long-range pixel interactions through multi-head attention and feed-forward networks, leveraging the relationship to model the restoration process.
@@
Intuitively, from the perspective of multi-modal complementarity, image fusion reveals the dominant information of different modalities, which in turn indicates the relative non-dominant information in the other modality, specifically tailoring regions that should be enhanced regardless of degradation issues. By leveraging the relative dominance in image fusion, the difficulty of image degradation recovery can be mitigated. Simultaneously, the use of image degradation tasks can further enhance the adaptability of image fusion to various degradation scenarios.
Based on this motivation, we propose a Dynamic Relative EnhAnceMent framework for Image Fusion (Dream-IF), jointly boosting the performance of image fusion and enhancement. Unlike most existing methods, Dream-IF first introduces the natural multi-modal complementarity from image fusion to enhance cross-modal image quality, especially for the corresponding non-dominant regions of the respective modality. To perform relative dominance-aware enhancement, we designed the relative enhancement (RE) module to dynamically capture the relative dominance of each modality, which is subsequently employed to facilitate cross-modal enhancement of the relatively weak-quality regions. Consequently, we can obtain the enhanced multi-modal features. By revealing the comprehensive nature of image fusion and converting the dominant regions of one modality to guide the enhancement of the other modality, we enhance the model’s capacity for image fusion in complex conditions, making the fusion model more robust to degradations. To the best of our knowledge, this work is the first to explore the intrinsic correlation of image enhancement and fusion using comprehensive relative dominance, rather than treating them as separate data-driven tasks or merely integrating them into a unified framework. Our model copes with robust image fusion for image degradation and even broader image enhancement. The main contributions can be summarized as follows:
  • We for the first time investigate the complementarity of image fusion from the perspective of image enhancement. Based on this, we propose a Dynamic Relative EnhAnceMent framework for Image Fusion, termed Dream-IF, qualified to perform robust image fusion for image degradation and even broader enhancement.
@@
Figure 2: An overview of the proposed Dream-IF. We introduce the Relative Enhancement block, which implicitly enhances non-dominant representations by leveraging relative information during the fusion and restoration process. This module captures the relative dominance inherent in the complementary nature of the image fusion model and uses it to facilitate both cross and self enhancement, ultimately producing the enhanced feature.
Related work
Image Fusion. Image fusion focuses on extracting effective information from multi-source images and producing an image containing complementary information. Traditional methods often perform image fusion by mathematical transformations, with multiscale analysis (Zhang and Blum 1999), sparse representation (Zhang et al. 2018), subspace learning (Fu et al. 2008), etc. With the development of deep learning, CNN-based methods have made significant advancements. DenseFuse (Li and Wu 2018) pioneered the use of deep learning models for fusing infrared and visible images. Inspired by transform-domain image fusion, IFCNN (Zhang et al. 2020) introduced a general image fusion framework for various fusion tasks. Furthermore, U2Fusion (Xu et al. 2020) addresses various fusion tasks through a unified, unsupervised image fusion network. Similar fusion tasks with aligned objectives can facilitate the integration of complementary information through cross-task commonality. More recently, TC-MoA (Zhu et al. 2024) dynamically selects task-customized mixture of adapters to capture task generality while preserving unique task characteristics. TTD (Cao et al. 2025) proposed an effective test-time fusion paradigm based on generalization theory proving dynamic fusion is superior to static fusion. Considering the potential low-quality of different sensors, some researchers (Deng and Dragotti 2020; Deng et al. 2023) have begun to recognize the correlation between image restoration and fusion. DIVFusion (Tang et al. 2023) enhances low-light images to produce a daytime-like fused image with improved visual perception. Different from these techniques, we first attempt to introduce the naturally complementary dominance of image fusion to dynamically enhance the image quality of non-dominant modality, improving the overall fusion results.
Image Restoration. In recent years, image restoration techniques (Liang et al. 2021; Zamir et al. 2022; Chen et al. 2022; Wang et al. 2022) have achieved significant progress. Particularly, blind restoration (Wu, Dong, and Qiao 2022; Huang and Xia 2020; Soh and Cho 2022) has gained considerable attention. Blind degradation problem is inverse where the degradation is not explicitly known, i.e., non-blind methods can model degradation using prior knowledge (e.g., a predefined blur convolution kernel) while blind methods only assume knowledge of the type of degradation (e.g., blurring)(Chihaoui, Lemkhenter, and Favaro 2024). ECycleGAN  (Wu, Dong, and Qiao 2022) tackled the blind reserve problem as an image-to-image task while preserving the fidelity of the reconstructed image. BlindDPS (Chung et al. 2023) solved blind inverse problems by constructing a separate diffusion prior for the forward operator. AirNet (Li et al. 2022) leveraged consistency to learn the degradation representation, recovering various degraded conditions. PromptIR (Potlapalli et al. 2024) introduces a prompt-based learning approach to encode degradation-specific features with prompts and dynamically guide all-in-one image restoration. Despite the extensive exploration of image restoration, these approaches fail to consider the inherent correlation between image restoration and fusion, specifically using the complementarity in image fusion to assist restoration.
Method
@@
We present the pipeline of our Dream-IF in Fig. 2, a dynamic relative enhancement framework acting on the fusion model, which leverages relative dominant prompts to perform cross-modal enhancement for non-dominant regions, thereby facilitating a more robust fusion.
The Overview of Dream-IF
@@
Dream-IF consists of encoders and decoders containing Restormer blocks (Zamir et al. 2022) corresponding to different modalities, which can be defined as $\mathcal{E}$ and $\mathcal{D}$ represent the encoder and decoder, respectively. Given the visible image $I^{v}\in\mathbb{R}^{H\times\ W\times 3}$ and the infrared image $I^{i}\in\mathbb{R}^{H\times\ W\times 3}$, the feature of each modality can be obtained as follows:
$F^{m}$ $=\mathcal{E}(I^{m}),$ (1)
$F_{i}^{m}$ $=\mathcal{D}_{i}(F_{i-1}^{m}),$ (2)
$I^{f}$ $=\mathcal{F}(\sum\nolimits_{k\in m}{F^{k}_{N}})$ (3)
@@
where $F^{m}\in\mathbb{R}^{h\times w}(m\in\{i,v\})$ represents the latent feature of the given modality, the $F^{m}_{i}$ denote the feature at the $i$th and $N$ is the number of decoder layer. Our goal is to fuse the complementary information from the two source images and obtain a fused image $I^{f}$. Thus, the fusion module can be formulated as $\mathcal{F}$, which consists of a Restormer block and a convolution layer to reconstruct the fused image from features.
Relative Enhancement
@@
The encoder and decoder tend to extract contemporary information from source images, which means that when a region of one modality is advantageous, another modality is probably disadvantageous. Inspired by the comprehensive nature of image fusion, we proposed a relative enhancement module to enhance the deficient regions of one modality by capturing the relative dominance of another modality, thus generating a more robust image enhancement and fusion.
Relative Dominance. As discussed above, the image fusion process inherently aims to exploit complementary information from multiple sources. To further enhance this capability, we introduce a mechanism that dynamically preserves complementary features while suppressing redundant information. Specifically, we define the fusion weight for each source feature as its Relative Dominance (RD), formulated as $\text{RD}^{m}=\sigma\text{Conv}(F^{m})$, where $\sigma$ denotes an activation function. The optimization objective is defined as:
$\min\mathbb{E}[\log(1-\sum_{h,w}^{N}\text{RD}^{(m)})].$ (4)
The fusion process aims at capturing the dominance of each source and leveraging it to enhance non-dominant areas. The visualizations of RD are shown in the Appendix.
Figure 3: Qualitative comparisons of various methods on representative images selected from the LLVIP, MSRS and M3FD datasets.
@@
Relative Enhancement. In image fusion, the dominance of one modality can be approximated by the non-dominance of another modality, which should be appropriately enhanced. Thus, the primary goal of the relative enhancement (RE) module is to boost the representation of non-dominant features by facilitating the relative dominance of another modality.
@@
To naturally integrate the RD and the feature to be strengthened, we present the cross enhancement (CE) and self enhancement (SE) module to enhance. The RD denotes the relative dominance from the other modality and indicates its deficient parts to be enhanced, thereby efficiently guiding the reconstruction process. Specifically, the CE module addresses deficiencies in one modality by complementing it with information from the other modality. It leverages the complementary nature of the modalities, filling in the gaps of the defective feature with the more dominant or reliable feature from the other modality. This ensures a more complete and accurate representation of the scene. The restoration process is guided by the dominant region in the other modality, where the strength of one modality helps identify areas in the other that require enhancement. By focusing on these dominant regions, the module effectively restores the missing or degraded information in the subordinate modality, resulting in a more coherent and refined output. It can be formulated as
@@
Further to enhance itself and restore degeneration, to make our model fundamentally resistant to degradation (Potlapalli et al. 2024), we generate a Self Enhancement (SE). This is calculated as:
$\hat{F}^{m}=F^{m}+\textbf{SE}(\textbf{RDP}(F^{m},RD,z),F^{m})$ (6)
where $z$ is a set of learnable parameters. The Relative Dominance Prompt (RDP) serves as an adaptive enhancement module, which generates restoration prompts dynamically based on the input feature $F^{m}$ and its corresponding relative dominance RD. Further architectural details are provided in the Appendix. Afterward, the refined feature is sent to the next block in the decoder, where it is transformed into the denoised and restored feature:
@@
[…]
Loss Function
@@
Moreover, to explicitly optimize the relative dominance (RD), we define the following loss function:
$\mathcal{L}_{RD}=\sum\nolimits_{j}^{h}\sum\nolimits_{k}^{w}\sum\nolimits_{m}^{N}RD^{m}_{i,j}-1.$ (9)
The overall loss function is defined as $\mathcal{L}=\mathcal{L}_{f}+\mathcal{L}_{\text{RD}}$.
Experiments
@@
Experimental Setting
Implementation Details. Our experimental framework is implemented using PyTorch (Paszke et al. 2019) and executed on an NVIDIA 3090 GPU. The proposed architecture adopts a U-Net-based design with a symmetric 4-level encoder-decoder structure, where the encoder extracts hierarchical features and the decoder reconstructs the output image. Each level of the encoder and decoder incorporates multiple Restormer (Zamir et al. 2022) blocks, with the number of blocks progressively increasing from higher to lower levels. The complete architectural details, including layer configurations and hyperparameters, are provided in the supplementary material for reproducibility.
@@
AG SF $Q^{abf}$ SSIM VIFF AG SF $Q^{abf}$ SSIM VIFF AG SF $Q^{abf}$ SSIM VIFF
TarDAL $4.921$ $18.207$ $0.410$ $1.080$ $0.537$ $1.914$ $5.944$ $0.170$ $0.930$ $0.380$ $4.331$ $15.766$ $0.292$ $0.731$ $0.486$
CDDFuse $5.403$ $18.495$ $0.582$ $1.184$ $0.660$ $3.779$ $11.570$ $0.585$ $1.303$ $0.717$ $4.800$ $14.709$ $0.521$ $1.379$ $0.578$
EMMA $5.560$ $17.190$ $0.552$ $1.192$ $0.648$ $3.775$ $11.559$ $0.621$ $1.324$ $0.738$ $\mathbf{5.362}$ $15.304$ $0.590$ $\underline{1.377}$ $\mathbf{0.711}$
Text-IF $\underline{5.682}$ $17.718$ $\underline{0.640}$ $1.074$ $\mathbf{0.813}$ $3.801$ $\underline{11.868}$ $\underline{0.624}$ $1.117$ $\mathbf{0.867}$ $5.034$ $15.494$ $\mathbf{0.644}$ $1.339$ $0.637$
TTD $5.509$ $\underline{19.914}$ $0.627$ $\underline{1.196}$ $0.730$ $3.705$ $11.527$ $0.549$ $\underline{1.349}$ $0.656$ $5.022$ $15.254$ $0.519$ $1.367$ $0.607$
Text-Dif $4.848$ $15.528$ $0.403$ $1.045$ $0.524$ $\underline{3.842}$ $11.506$ $0.436$ $0.935$ $0.703$ $2.810$ $8.465$ $0.149$ $1.122$ $0.165$
FusionBooster $5.476$ $15.464$ $0.407$ $0.978$ $0.551$ $3.208$ $9.012$ $0.420$ $1.043$ $0.630$ $3.559$ $10.245$ $0.396$ $1.297$ $0.515$
Dream-IF $\mathbf{5.926}$ $\mathbf{19.992}$ $\mathbf{0.679}$ $\mathbf{1.198}$ $\underline{0.800}$ $\mathbf{3.962}$ $\mathbf{12.293}$ $\mathbf{0.631}$ $\mathbf{1.362}$ $\underline{0.849}$ $\underline{5.124}$ $\mathbf{16.198}$ $\underline{0.614}$ $\mathbf{1.387}$ $\underline{0.704}$
Table 1: Quantitative comparison with SOTAs without degradation. The bold/underline indicates the best and runner-up.
Datasets. Our experiments are conducted on three widely recognized publicly available datasets: LLVIP, MSRS, and M3FD. Further details about the datasets are provided in the supplement. Following (Zhu et al. 2024), utilizing $12,025$ image pairs from the LLVIP dataset with random degradation for training and $70$ image pairs for testing. To rigorously assess the generalization capability of the proposed method, we tested the model on both the MSRS and M3FD datasets without any fine-tuning.
Competing Methods. We compared our approach with seven recent competing methods, including TarDAL (Liu et al. 2022), CDDFuse (Zhao et al. 2023), EMMA (Zhao et al. 2024), Text-IF (Yi et al. 2024), TTD (Cao et al. 2025), Text-DiFuse (Zhang, Cao, and Ma 2025), and FusionBooster (Cheng et al. 2025).
Evaluation Metrics. We evaluated the fusion results quantitatively on five metrics, including the average gradient (AG), spatial frequency (SF), gradient-based similarity measurement $Q^{abf}$ (Piella and Heijmans 2003), structural similarity (SSIM) (Wang et al. 2004), and visual information fidelity for fusion (VIFF) (Han et al. 2013).
Comparison without Degradation
@@
Qualitative Comparison. Our comprehensive evaluation reveals that, while all competing fusion methods are capable of combining primary features from infrared and visible images to some extent, the proposed method demonstrates clear advantages that significantly enhance fusion quality. Our method excels in emphasizing the critical characteristics of infrared images while preserving complementary visible information. As illustrated in Fig. 3 (LLVIP dataset, red boxes), the visible images in low-light conditions are often blurred, whereas the infrared images provide essential complementary details. Competing methods, however, suffer from significant infrared content loss in the fusion results, as evidenced by the dark and indistinct details of the cars in the green boxes. Although CDDFuse shows competitive visual performance in the red boxes, it fails to retain sufficient infrared information, leading to suboptimal fusion in the green boxes. In contrast, our method enhances vulnerable modalities through a robust feature integration mechanism, effectively preserving both infrared and visible information. This results in superior perceptual quality and improved information retention, as demonstrated by the clear and detailed fusion results in both highlighted regions. The robust experimental results show consistent performance on both the MSRS and M3FD datasets.